OpenStudy (anonymous):

I have to graph the line using this equation; 3x+2y=-4 PLEASE HELP? =)

OpenStudy (anonymous):

First of all, solve for y. y=???

OpenStudy (phi):

You could try to find 2 points that are on the line. Then plot the 2 points and draw a line through the points (connect the dots...) one way to find 2 points is replace x with some nice easy number and solve for y for example, if x is -1 you would do this 3x+2y=-4 3 * -1 + 2y = -4 (we replaced x with -1) that simplifies to -3 + 2y = -4 add +3 to both sides -3+3 + 2y = -4 + 3 simplify some more: 2y = -1 divide both sides by 2 y = -1/2 that means that (-1, -1/2) is on the line now find another point. Try using x=0 and solve for y

OpenStudy (anonymous):

(-4/3, 0) , ( 0,-2) i think these are my two points soemtimes i get it sometimes i don't i think my main issue is how after i submitted my answer for another problem the explanation they gave me went from 4x+3y+6 =0 to 3y=-4x-6 to y=4/3x -2 and i don't understand how they were working it out

OpenStudy (phi):

yes, re-writing the equation in the form \[ y= \frac{4}{3}x -2 \] is a good way. This tells you the slope is 4/3 (up 4 and over three) You use algebra to "solve for y" It might be easiest to explain if you look at this http://www.khanacademy.org/math/algebra/linear-equations-and-inequalitie/equation-of-a-line/v/converting-to-slope-intercept-form

OpenStudy (anonymous):

ohh yeahh i used to use knan i forgot about that site, thanks =)